OpenStudy (anonymous):

(5x - 7y + 8z) - (-2x + 3y -z)

OpenStudy (callisto):

(5x - 7y + 8z) - (-2x + 3y -z) = 5x - 7y + 8z + 2x - 3y +z = 5x + 2x -7y -3y +8z + z = combine the like terms and get the answer!